The perimeter of a rectangle is 54m. The length is 6 m more than twice the width. What is the length and what is the width? Please advise. Thank you.
Answer by checkley77(12569)

54=2L+2W
L=2W+6
54=2(2W+6)+2W
54=4W+12+2W
54=6W+12
6W=54-12
6W=42
W=42/6
W=7 ANS.
L=2*7+6
L=14+6
L=20 ANS.
PROOF:
54=2*20+2*7
54=40+14
54=54